Required information

[The following information applies to the questions displayed below.]

You have just been hired as a financial analyst for Lydex Company, a manufacturer of safety helmets. Your boss has asked you to perform a comprehensive analysis of the company’s financial statements, including comparing Lydex’s performance to its major competitors. The company’s financial statements for the last two years are as follows:

Lydex Company
Comparative Balance Sheet
This Year Last Year
Assets
Current assets:
Cash $ 970,000 $ 1,210,000
Marketable securities 0 300,000
Accounts receivable, net 2,740,000 1,840,000
Inventory 3,610,000 2,100,000
Prepaid expenses 260,000 200,000
Total current assets 7,580,000 5,650,000
Plant and equipment, net 9,540,000 9,060,000
Total assets $ 17,120,000 $ 14,710,000
Liabilities and Stockholders' Equity
Liabilities:
Current liabilities $ 4,020,000 $ 3,000,000
Note payable, 10% 3,680,000 3,080,000
Total liabilities 7,700,000 6,080,000
Stockholders' equity:
Common stock, $70 par value 7,000,000 7,000,000
Retained earnings 2,420,000 1,630,000
Total stockholders' equity 9,420,000 8,630,000
Total liabilities and stockholders' equity $ 17,120,000 $ 14,710,000
Lydex Company
Comparative Income Statement and Reconciliation
This Year Last Year
Sales (all on account) $ 15,870,000 $ 13,680,000
Cost of goods sold 12,696,000 10,260,000
Gross margin 3,174,000 3,420,000
Selling and administrative expenses 1,106,000 1,608,000
Net operating income 2,068,000 1,812,000
Interest expense 368,000 308,000
Net income before taxes 1,700,000 1,504,000
Income taxes (30%) 510,000 451,200
Net income 1,190,000 1,052,800
Common dividends 400,000 526,400
Net income retained 790,000 526,400
Beginning retained earnings 1,630,000 1,103,600
Ending retained earnings $ 2,420,000 $ 1,630,000

To begin your assignment you gather the following financial data and ratios that are typical of companies in Lydex Company’s industry:

Current ratio 2.4
Acid-test ratio 1.1
Average collection period 40 days
Average sale period 60 days
Return on assets 9.4 %
Debt-to-equity ratio 0.7
Times interest earned ratio 5.7
Price-earnings ratio 10

2. You decide next to assess the company’s stock market performance. Assume that Lydex’s stock price at the end of this year is $96 per share and that at the end of last year it was $64. For both this year and last year, compute: (Round your "Percentage" answers to 1 decimal place and other intermediate and final answers to 2 decimal places.)

a. The earnings per share.

b. The dividend yield ratio.

c. The dividend payout ratio.

d. The price-earnings ratio.

e. The book value per share of common stock.
Answer #1
a) Earning Per Shares = Net Income/ Average Number Of Shares Outstanding
= $1190000/100000 = $1052800/100000
= $ 11.9 = $ 10.53
b) The dividend yield ratio = dividend per share/Market price per share
This Year Last Year
Dividend per share =$400000/100000 =$526400/100000
=$4 per share =$5.26 per share
The dividend yield ratio. =$4/96 =$5.26/64
=4.2% =8.2%
c)
Payout Ratio = Dividend Per Share/ Earning Per Share
This Year Last Year
= $4/11.9 = $5.26/10.53
= 33.6 % =50 %
d) Price Earning Ratio = Market price per share/ Earning per share
This Year Last Year
=$96/11.9 =$64/10.53
=8.07 =6.08
e) Book Value per share = Common stock holders equity/ number of common shares
This Year Last Year
=9,420,000/100000 =8,630,000/100000
=$94.20 =$86.30
Let me know the wrong answer, if any. I will correct it immediately. Please upvote.